an=3an-1-4,a0=1


1
Expert's answer
2021-02-25T05:08:35-0500

"a_n=a_{n-1}-4,a_0=1\\
a_1=a_0-4=1-4=-3\\
a_2=a_1-4=(1-4)-4\\
a_3=a_2-4=(1-4-4)-4\\
\vdots\\
a_n=1-4n"